OpenStudy (anonymous):

Help me subtract this problem. [(4x)/2x-7]-[(x-7)/2x-7]

OpenStudy (anonymous):

I put my final answer as \[\frac{ 3x-7 }{ 2x-7}\] but I got it wrong because it says that the answer is \[\frac{ 3x+7 }{ 2x-7}\] Why did I need to change the sign of 7?

OpenStudy (anonymous):

@ashley_f97

OpenStudy (phi):

\[ \frac{4x}{2x-7} - \frac{(x-7)}{2x-7}\] both fractions have the same denominator, so you can subtract the numerators \[ \frac{4x-(x-7)}{2x-7} \] now notice you are doing 4x- (x -7) the - subtracts *everything inside the parens* it is the same as 4x -x - (-7) the -(-7) becomes +7 (two minuses make a plus. the top is 4x - x +7 or 3x+7

OpenStudy (phi):

another way to do this is 4x - (x-7) is the same as 4x + -1*(x-7) now "distribute" the -1, by multiplying both terms inside the parens by -1 4x + -1*x + -1*-7 simplify to 4x -x +7 and then 3x+7

OpenStudy (anonymous):

Thank you!

OpenStudy (phi):

yw